NEW ASSIGNMENTS IN THE NEAR INFRARED SPECTRUM OF METHYLENE

ANDREW J. MARR and \underline{TREVOR J. SEARS}, Department of Chemistry, Brookhaven National Laboratory, Upton, New York 11973-5000; BOR-CHEN CHANG, Department of Chemistry, National Central University, Chung-Li, Taiwan 32054.

Transitions to five previously unobserved levels, have been identified in a newly recorded region (10 000 cm^{−1} - 10 600 cm^{−1}) of the near infrared spectrum of the b̃^{1}B_{1} ← ã^{1}A_{1} system of CH_{2}. Four states were observed via transitions from the ã(0,0,0) state: K = 1 ã(0,9,0), K = 2 b̃(0,1,0), K = 2 ã(1,6,0), and K = 3 b̃(0,1,0), and one via hot-band transitions from the ã(0,1,0) state: K = 3 ã(0,10,0). The observation of levels with predominantly ã state character is dependent on strong Renner-Teller coupling of the two singlet states which correlate to a ^{1}Δ_{g} state at linearity. Hot band transitions to two other states previously identified at shorter wavelengths are observed. Comparison with recent theoretical calculations^{a}^{,}^{b}^{,}^{c} shows good agreement in most cases. These assignments still account for only approximately 20% of the lines observed in this region; speculations on the identity of the unassigned transitions will be discussed. Acknowledgements: The work was carried out at Brookhaven National Laboratory under Contract No. DE-AC02- 76CH00016 with the U.S. Department of Energy and supported by its Division of Chemical Sciences, Office of Basic Energy Sciences. BCC also thanks the National Science Council, Taiwan for its support of this work under Grant No. NSC87-2113-M-008-0009.

aP. Jensen, M. Brumm, W. P. Kraemer, and P. R. Bunker, J. Mol. Spectrosc. \underline{171}, 31 (1995).

bW. H. Green, Jr., N. C. Handy, P. J. Knowles, and S. Carter, J. Chem. Phys. \underline{94}, 118 (1991).

cG. Duxbury, A. Alijah, B. D. McDonald, and Ch. Jungen, J. Chem. Phys. \underline{108} , 2351 (1998).